OpenStudy (anonymous):

Multiply. Try to recognize the type of product before multiplying (4n+7)(4n-7)

OpenStudy (diyadiya):

This is of the form (a+b)(a-b) (a+b)(a-b)=a^2-b^2

OpenStudy (anonymous):

i dont understand

OpenStudy (diyadiya):

Ok don't use that Multiply each terms 4n(4n-7)+7(4n-7)

OpenStudy (anonymous):

so it would b 16n what do i do with the 7 sorry math is very hard for me

OpenStudy (diyadiya):

4n(4n-7) + 7(4n-7) 16n-7*4n+7*4n-7*7

OpenStudy (anonymous):

so now do i combine like terms

OpenStudy (diyadiya):

Sorry its actually 16n^2 (4n*4n=16n^2) yeah now combine like terms check this too ,this is how we got it